The reaction of carbon dioxide(g) with hydrogen(g) to form acetylene(g) (C2H2) and water(g) proceeds as follows:
2 CO2(g) + 5 H2(g) C2H2(g) + 4 H2O(g)
When 15.8 grams of CO2(g) react with sufficient H2(g) , 8.35 kJ of energy are absorbed .

What is the value of H for the chemical equation given?

ΔHrxn = ________ kJ
1
Expert's answer
2019-04-08T07:56:39-0400
"n=\frac{m}{M}"

"n(CO_2)=\frac{15.8g}{44\frac{g}{mol}}=0.359 mol"

When 0.359 mol of "CO_2" react with sufficient amount of "H_2" 8.35 kJ of energy are absorbed.

When 2 moles of "CO_2" react with sufficient amount of "H_2" xkJ of enerhy is absorbed.

Solve the proportion:


"\frac{0.359 mol}{2 mol} = \frac {8.35 kJ}{x}"

"x= 46.52 kJ"